. Find the amount owing on a loan of $7400 at the end of 3 years, with interest at 10.9%/annum, compounded monthly.

Answer by ewatrrr(10682)


Hi
Note: A= p(1-r/n) ^nt
amount owing on a loan of 7400,after 3 years: t = 3
with annual interest=0.109 compounded monthly: n =12 r/n = .009
amount owing =7400*(1-0.109/12)^(12*3)= $5328.05